It is estimated that 9% of those taking the quantitative methods portion of the certified public accountant (CPA) examination fail that section. Sixty three students are taking the examination this Saturday

. a-1. How many would you expect to fail? (Round the final answer to 2 decimal places.)

Number of students

a-2. What is the standard deviation? (Round the final answer to 2 decimal places.)

Standard deviation

b. What is the probability that exactly five students will fail? (Round the final answer to 4 decimal places.)

Probability

c. What is the probability at least five students will fail? (Round the final answer to 4 decimal places.)

Probability
Answer #1

a-1)

Number of students fail = Total students*Fail % = 63*0.09 = 5.67

a-2)

standard deviation = sqrt(Total students*Fail %*Pass %) = sqrt(63*0.09*0.91) = 2.27

b)

P(x=T) = Total studentsCFail*(Fail %)^Number of Fail*(Pass%)^Number of pass

P(x=5) = 63C5*((0.09)^5)*((0.91)^58) = 0.1748